Find the volume of the largest sphere that can be inscribed in a cube of edge 12 cm.
Answer by stanbon(75887)

Find the volume of the largest sphere that
can be inscribed in a cube of edge 12 cm.
------
Diameter of the Sphere = 12 cm
Radius = 6 cm
-----
Sphere Volume:: V = (4/3)pi*r^3
----
Ans: V = (4/3)pi*6^3 = 904.78 cm^3
--------------
Cheers,
Stan H.
-------------